### v1

# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

Abstract
This study considers the estimation of the average treatment effect (ATE). For ATE estimation, we estimate the propensity score through direct bias-correction term estimation. Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ is an outcome. In ATE estimation, the bias-correction term $h_{0}(X_{i},D_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ plays an important role, where $e_{0}(X_{i})$ is the propensity score, the probability of being assigned treatment $1$. In this study, we propose estimating $h_{0}$ (or equivalently $e_{0}$) by directly minimizing the prediction error for $h_{0}$ without knowing $h_{0}$ in advance. After showing a basic result with least squares, we present a general framework for this direct bias-correction term estimation approach from the perspective of Bregman divergence minimization, which also generalizes the Riesz regression and covariate balancing.
1 Introduction
We consider the problem of estimating the average treatment effect (ATE) in causal inference (Imbens & Rubin, 2015). Methods for estimating ATEs are typically designed to eliminate bias arising from treatment assignment and the estimation of nuisance parameters, aiming for (asymptotic) unbiasedness and efficiency.
In ATE estimation, the propensity score, the probability of treatment assignment given covariates, plays a crucial role (Rosenbaum & Rubin, 1983). For example, the inverse probability weighting (IPW) estimator, also known as the Horvitz-Thompson estimator (Horvitz & Thompson, 1952), estimates the ATE by computing a weighted average of outcomes using the inverse of the propensity score. In the one-step bias correction method (van der Vaart, 2002), a bias-correction term depending on the propensity score is added to an initial estimator.
In this study, we focus on the role of the propensity score as a bias-correction component and propose a novel method to directly estimate either the propensity score or the bias-correction term itself. The bias-correction term is fundamental to IPW-type estimators and the one-step bias correction method, and its accurate estimation can substantially improve ATE estimation. Importantly, our objective is not to estimate the propensity score, but rather the bias-correction term, in which the propensity score appears inversely and is weighted by the treatment indicator. Since estimating the propensity score itself is not the target, we hypothesize that directly estimating the bias-correction term will lead to improved performance. To this end, we propose estimating the bias-correction term by minimizing the empirical risk targeted for the true bias-correction term.
The technical challenge is that the target variable remains unobserved, even when we directly estimate the bias-correction term or the propensity score. To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012). In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown. It is known that empirical risk minimization is feasible even without knowledge of the true propensity score. Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting.
Our motivation is closely aligned with studies on covariate balancing (Imai & Ratkovic, 2013) and Riesz regression (Chernozhukov et al., 2022a), which also aim to improve ATE estimation by appropriately estimating the propensity score or the bias-correction term. Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property. Chernozhukov et al. (2022a) proposes Riesz regression which represents the bias-correction term as the Riesz representer. Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al. (2022a) by using the DRE techniques. Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al. (2011). From this generalization, we connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing (Chan et al., 2015).
1.1 ATE estimators and bias correction
We begin by formulating the problem. There are two treatments, denoted by $1$ and 0.111In some cases, only treatment $1$ is referred to as the treatment, while treatment 0 is referred to as the control. For simplicity, we refer to them as treatment $1$ and treatment 0 throughout this study. For each treatment $d\in\{1,0\}$, let $Y(d)\in{\mathbb{R}}$ denote the potential outcome under treatment $d$. The treatment assignment indicator is denoted by $D\in\{1,0\}$, and the observed outcome is given by $Y=\mathbbm{1}[D=1]Y(1)+\mathbbm{1}[D=0]Y(0)$, meaning that we observe $Y(d)$ only if the unit is actually assigned to treatment $d$. Each unit is characterized by $K$-dimensional covariates $X\in\mathcal{X}\subset{\mathbb{R}}^{K}$, where $\mathcal{X}$ denotes the covariate space. For $n$ units indexed by $1,2,\dots,n$, let $\mathcal{D}\coloneqq\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ denote the observed data, where each $(X_{i},D_{i},Y_{i})$ is an i.i.d. copy of $(X,D,Y)$ generated from an underlying distribution $P_{0}$. Our goal is to estimate the ATE, defined as
$\tau_{0}\coloneqq{\mathbb{E}}\big[Y(1)-Y(0)\big],$
where the expectation is taken over the distribution $P_{0}$. Note that we can also apply our method for the ATE for the treated group (ATT). For the details about ATT estimation, see Appendix C.
Let $e_{0}(X)=P_{0}(D=1\mid X)$ denote the probability of assigning treatment $1$ given covariates $X$, which is known as the propensity score. Throughout this study, we impose the following condition, commonly referred to as the overlap assumption:
Assumption 1.1.
There exists a constant $C>0$ independent of $n$ such that $C<e_{0}(x)<1-C$ for all $x\in\mathcal{X}$.
When $e_{0}(x)$ is not constant, a distributional shift arises between the observed outcomes in the treatment and control groups, denoted by $\mathcal{G}_{1}$ and $\mathcal{G}_{0}$, respectively, where $\mathcal{G}_{d}\coloneqq\{i\in\{1,2,\dots,n\}\colon D_{i}=d\}$. This shift induces bias in the sample mean, $\frac{1}{|\mathcal{G}_{d}|}\sum_{i\in\mathcal{G}_{d}}Y_{i}=\frac{1}{|\mathcal{G}_{d}|}\sum_{i\in\mathcal{G}_{d}}Y_{i}(d)$, which deviates from ${\mathbb{E}}\left[Y(d)\right]$ and thus prevents the sample mean difference, $\frac{1}{|\mathcal{G}_{1}|}\sum_{i\in\mathcal{G}_{1}}Y_{i}-\frac{1}{|\mathcal{G}_{0}|}\sum_{i\in\mathcal{G}_{0}}Y_{i}$, from being an unbiased estimator of the ATE.
To address this issue, several debiased estimators have been proposed under standard regularity conditions. In this section, we introduce two representative estimators, the inverse probability weighting (IPW) estimator and the augmented IPW (AIPW) estimator, as follows:
IPW estimator.
$\widetilde{\tau}^{\mathrm{IPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]Y_{i}}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]Y_{i}}{1-e_{0}(X_{i})}\right)=\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}\right)Y_{i}$.
AIPW estimator.
$\widetilde{\tau}^{\mathrm{AIPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\left(\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}\right)\left(Y_{i}-\mu_{0}(D_{i},X_{i})\right)+\mu_{0}(1,X_{i})-\mu_{0}(0,X_{i})\right)$, where $\mu_{0}(d,X)$ is the expected conditional outcome ${\mathbb{E}}\left[Y(d)\mid X\right]$ of treatment $d$ given $X$. The AIPW estimator is also known as the doubly robust (DR) estimator (Bang & Robins, 2005).
Bias-correction term. In both estimators, the term
$h_{0}(D,X)\coloneqq h(D,X;1/e_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
is crucial. This term, referred to as the bias-correction term, is central to ATE estimation (Schuler & van der Laan, 2024). A common approach is to estimate $e_{0}$ using logistic regression and then plug the resulting estimate $\widehat{e}^{{\mathrm{L}}}_{n}$ into $h$. Note that in automatic debiased machine learning, the term is also referred to as the Riesz representer (Chernozhukov et al., 2022b).
For example, in a typical one-step bias correction, we first construct an ATE estimator as $\widehat{\tau}^{\mathrm{DM}}_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\mu}_{n}(1,X)-\widehat{\mu}_{n}(0,X)\right)$, where $\widehat{\mu}_{n}$ is an estimator of $\mu_{0}$. This estimator is known as the direct method (DM) or naive plug-in estimator. To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(X_{i},D_{i})$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
In this process, estimating the propensity score $e_{0}$ itself is a challenging problem, independent of ATE estimation. The well-known Vapnik principle states that “when solving a problem of interest, do not solve a more general problem as an intermediate step” (Vapnik, 1998). Following this principle, this study aims to estimate $h_{0}(D,X)$, or equivalently $e_{0}$, by directly minimizing the estimation error of $h_{0}(D,X)$. We emphasize that while $e_{0}$ is estimated in our approach, the primary estimation target is not $e_{0}$ but $h_{0}$.
1.2 Our contributions
Our first contribution is the proposal of a framework for direct bias-correction term estimation. We estimate the bias-correction term by directly minimizing the estimation error for the true bias-correction term $h_{0}$. To model the bias-correction term $h_{0}$, we define the inverse propensity score as $r_{0}(1,X)=\frac{1}{e_{0}(X)}$ and $r_{0}(0,X)=\frac{1}{1-e_{0}(X)}$. Let $\mathcal{R}$ be a model (hypothesis class) for $r_{0}$, and consider approximating $r_{0}$ using some $r\in\mathcal{R}$. We approximate $r_{0}$ by the minimizer $r^{*}\in\mathcal{R}$ of the mean squared error (MSE) ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$. Here, we estimate $h_{0}(D,X)=h(D,X;r_{0})$ using a model $h(D,X;r)$ that depends on $r$. We emphasize again that although we bypass the estimation of $r_{0}$, our primary estimation target is $h_{0}$, not $r_{0}$.
Since this expected squared error involves the unknown function $r_{0}$, direct optimization is infeasible. However, we show that minimizing this expected squared error with respect to $r$ is equivalent to minimizing ${\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$, which does not depend on the unknown function. That is, we establish the equivalence: $r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. The obtained squared error can then be approximated using a sample-based empirical risk function.
Our second main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation. Since we estimate $r_{0}$ using empirical risk minimization, we establish bounds on the estimation error using empirical process theory. Furthermore, we present examples of ATE estimators that incorporate the bias-correction term estimated using our framework and conduct simulation studies. Using standard ATE estimation techniques, we demonstrate that our method yields a $\sqrt{n}$-consistent ATE estimator.
Our third main contribution is the generalization of our framework. From the perspective of Bregman divergence minimization, we extend our framework to provide a more general methodology for direct bias-correction term estimation. Under this framework, various estimation strategies can be incorporated to enhance bias-correction term estimation.
Our fourth contribution is the unification of the existing literature: Riesz regression and covariate balancing. If we use the squared loss function in the Bregman divergence minimization, we obtain the same loss as the one used in Riesz regression. If we use a different function in the Bregman divergence, we recover a special case of the tailored loss for covariate balancing proposed in Zhao (2019), which is also equivalent to empirical balancing (Chan et al., 2015; Hainmueller, 2012).
2 Direct bias-correction term estimation
In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the MSE ${\mathbb{E}}\left[\big(h_{0}(D,X)-h(D,X)\big)^{2}\right]$ over $h:\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$. Here, since $h_{0}$ contains the unknown function $r_{0}$, this minimization problem is inherently intractable. However, we demonstrate that empirical risk minimization remains feasible without explicit knowledge of $r_{0}$.
2.1 Least squares
Rather than directly modeling $h_{0}$, we model either $r_{0}$ or $e_{0}$ within $h_{0}$. That is, we model $h_{0}$ via a model of $r_{0}$ as $h(D,X;r)$, while $h_{0}(D,X)$ is given as $h_{0}(D,X;r_{0})$. We then aim to estimate $r_{0}$ or $e_{0}$. Given a set $\mathcal{R}$ of functions $r\colon{\mathcal{X}}\to(1,\infty)$, we approximate the bias-correction term by solving
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right].$
If $r_{0}\in\mathcal{R}$, then it follows that $r_{0}(\cdot,\cdot)=r^{*}(\cdot)$ and $h(\cdot,\cdot;r^{*})=h_{0}(\cdot,\cdot)$. This MSE is associated with the MSE between the ATE $\tau_{0}$ and some estimator using the bias-correction term. Since $e_{0}(x)\in(0,1)$, a model $r(x)$ takes a value in $(1,\infty)$.
Example. For example, for $\widehat{\tau}^{\mathrm{IPW}}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}h(D_{i},X_{i};r)Y_{i}$, we have ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}+\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}\right)^{2}\right]+{\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$, where recall that we defined $\widetilde{\tau}^{\mathrm{IPW}}$ in the Introduction. Here, we have ${\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]=\frac{1}{n}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}{\mathbb{E}}\left[Y^{2}\mid D,X\right]\right]$. Thus, the MSE ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$ is closely connected to the MSE ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$.
Direct estimation without $r_{0}$. The essential problem in this approach is that the target $h_{0}$ or $r_{0}$ is unknown; therefore, this optimization looks infeasible. However, even if we do not know $h_{0}$, we can reformulate the optimization problem into an equivalent form that does not involve $h_{0}$ as
$r^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big].$
This reformulation is derived as follows: $\min_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\Big[h(D,X;r_{0})^{2}-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\Big]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. From the second to the third line, we omit the term $h_{0}(D,X)^{2}$ since it does not affect the optimization. From the third to the fourth line, we use the following for the first term: ${\mathbb{E}}\left[h(D,X;r_{0})h(D,X)\right]={\mathbb{E}}\Bigg[\left(\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}\right)\Big(\mathbbm{1}[D=1]r(1,X)-\mathbbm{1}[D=0]r(0,X)\Big)\Bigg]={\mathbb{E}}\left[\frac{\mathbbm{1}[D=1]}{e_{0}(X)}r(1,X)+\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[\frac{e_{0}(X)}{e_{0}(X)}r(1,X)+\frac{1-e_{0}(X)}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[r(1,X)+r(0,X)\right]$.
Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ (or the unknown $r_{0}$) itself. As discussed in the following subsection, this objective function can be easily approximated using observations.
2.2 Empirical risk minimization
We then estimate $r_{0}$ by solving the following empirical risk minimization:
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where the empirical risk is given as
$\widehat{{\mathcal{L}}}_{n}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r(1,X_{i})-2r(0,X_{i})+\mathbbm{1}[D_{i}=1]r(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r(0,X_{i})^{2}\Big),$
and $J\colon{\mathcal{R}}\to{\mathbb{R}}^{+}$ is a regularization term with a penalty coefficient $\lambda>0$.
This type of estimation method is referred to as least-squares importance fitting (LSIF, Kanamori et al., 2009) in the literature on density-ratio estimation (DRE).
3 Estimation error analysis
This section provides an estimation error analysis for $r_{0}$ estimated by the direct bias-correction term estimation method. For simplicity, throughout the analysis, we use the following model for $e_{0}$: $e(x)=\frac{1}{1+\exp(-f(x))}$, where $f\colon{\mathcal{X}}\to{\mathbb{R}}$ is a function belonging to a set $\mathcal{F}$}. We use various models for $\mathcal{F}$}, including linear models, RKHS, and neural networks. Then, the model of $r_{0}$ is given as $r(1,x)=1+\exp(-f(x)),\quad r(0,x)=\exp(f(x))$. Let $f_{0}$ be the true function for $e_{0}$; that is, $e_{0}(X)=\frac{1}{1+\exp(-f_{0}(x))}$.
3.1 Linear models
First, we consider the case where $f_{0}$ belongs to a linear model. There exists $\theta_{0}\in\Theta^{K}$ such that $e_{0}(x)=\frac{1}{1+\exp\big(-x^{\top}\theta_{0}\big)}$, where $\Theta$ is a compact parameter space. Under this assumption, we have $r_{0}(1,X)=1+\exp\big(-x^{\top}\theta_{0}\big)$ and $r_{0}(0,X)=1+\exp\big(x^{\top}\theta_{0}\big)$. Let us use $r_{\theta}(1,X)=1+\exp\big(-x^{\top}\theta\big)$ and $r_{\theta}(0,X)=1+\exp\big(x^{\top}\theta\big)$ as a model for $r_{0}$. Then, we define the estimator as $\widehat{r}^{\mathrm{Lin}}_{n}=r_{\widehat{\theta}_{n}}$, where $\widehat{\theta}_{n}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$. Here, we set $\lambda=0$.
We show the asymptotic normality of this estimator.
Assumption 3.1.
We assume the following: (i) $\theta_{0}$ is in the interior of $\Theta$; (ii) $\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$ is twice continuously differentiable on some neighborhood $\mathcal{M}$} of $\theta_{0}$ (with probability one); (iii) $\sqrt{n}\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})\xrightarrow{d}{\mathcal{N}}(0,\Sigma_{0})$; (iv) for any sequence $\tilde{\theta}_{n}\xrightarrow{p}\theta_{0}$, it holds that $\frac{\partial^{2}}{\partial\theta\partial\theta^{T}}\widehat{{\mathcal{L}}}_{n}(r_{\theta})-B_{0}\xrightarrow{p}0$, for some non-stochastic $K\times K$ matrix $B_{0}$ that is nonsingular.
Then, from the asymptotic theory of extremum estimators, we obtain the following result:
Lemma 3.2.
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,B_{0}^{-1}\Sigma_{0}B_{0}^{-1})$.
From the Taylor expansion of $r_{\theta}(x)$ around $\theta=\theta_{0}$, Lemma 3.2 immediately yields the following result:
Theorem $3.$
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{r}^{\mathrm{Lin}}_{n}(x)-r_{0}(x)\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,\kappa(x))$, where $\kappa(x)\coloneqq\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)^{\top}B_{0}^{-1}\Sigma_{0}B_{0}^{-1}\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)$.
3.2 RKHS
Next, we investigate the case with RKHS regression. Let ${\mathcal{R}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define the estimator as $\widehat{r}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda\|r\|^{2}_{{\mathcal{H}}}$, where $\|\cdot\|^{2}_{{\mathcal{H}}}$ is the RKHS norm. We analyze the estimation error by employing the results in Kanamori et al. (2012), which study RKHS-based LSIF in DRE.
We define the following localized class of RKHS functions as a technical device: ${\mathcal{R}}^{\text{RKHS}}_{M}\coloneqq\big\{r\in{\mathcal{R}}^{\text{RKHS}}\colon I(r)\leq M\big\}$. We then make the following assumption using this localized class.
Assumption 3.4.
There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{R}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
For the details of the definition of the bracketing entropy, see Appendix F and Definition 2.2 in van de Geer (2000).
Under these preparations, we establish an estimation error bound.
Theorem 3.5 ($L_{2}$-norm estimation error bound).
Suppose that Assumptions 1.1 and 3.4 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, then we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|\big(1-e_{0}(X)\big)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=O_{P_{0}}\left(\lambda^{1/2}\right)$.
The proof is provided in Appendix F, following the approach of Kanamori et al. (2012). The parameter $\gamma$ is determined by the function class to which $f_{0}$ belongs.
3.3 Neural networks
This section provides an estimation error analysis when we use neural networks for $\mathcal{R}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022).
Feedforward neural networks (FNNs). We define FNNs as follows:
Definition 3.6 (Feedforward neural networks. From Zheng et al. (2022)).
Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$. Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs $r_{\theta}(1,x)=1/e_{\theta}(x)$ and $r_{\theta}(0,x)=1/(1-e_{\theta}(x))$, where $e_{\theta}\colon{\mathbb{R}}^{K}\to{\mathbb{R}}$ with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022).
We model $r_{0}$ by $r(1,X)=\frac{1}{e(x)}$ and $r(0,X)=\frac{1}{1-e(x)}$, where $e(x)=\frac{1}{1+\exp(-f(x))}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let us denote such an estimator by $r_{f}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let $f_{0}$ denote the true function such that $r_{f_{0}}=1+\exp(-f_{0}(x))$ holds. Then, the estimator is denoted as $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(r_{f})$.
Estimation error analysis. For the estimator, we can prove an estimation error bound. Let us make the following assumption.
Assumption 3.7.
There exists a constant $0<M<\infty$ such that $\|f_{0}\|_{\infty}<M$, and $\|f\|_{\infty}\leq M$ for any $f\in{\mathcal{F}}^{\mathrm{FNN}}$.
Then, we prove the following estimation error bound:
Theorem 3.8 (Estimation error bound for neural networks).
Suppose that Assumption 3.7 holds. Also assume $r_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$. Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{FNN}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|(1-e_{0}(X))\left(\widehat{r}^{\mathrm{FNN}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
The proof is provided in Appendix G, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
4 Example about the AIPW estimator
This section introduces the AIPW estimator with nuisance parameters estimated using our proposed direct bias-correction term estimation. We prove that under certain conditions, the proposed estimator is asymptotically normal. Note that this result is well known in the literature except for the use of nuisance parameters estimated via our direct bias-correction term estimation. The purpose of this section is not to provide novel methodological or theoretical results but to present an application of our proposed method.
We analyze the AIPW estimator with an estimated propensity score. Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
We first make the following assumption.
Assumption 4.1 (Donsker condition).
The hypothesis classes $\mathcal{R}$} and $\mathcal{M}$} belong to the Donsker class.
For example, this assumption holds when the bracketing entropy is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if this assumption does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML Chernozhukov et al., 2018, ,). In DML, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$. For more details, see Chernozhukov et al. (2018).
Assumption 4.2 (Convergence rate).
$\big\|\widehat{r}-r_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{r}-r_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
Under these assumptions, we show the asymptotic normality of $\widetilde{\tau}^{\mathrm{AIPW}}_{n}$. We omit the proof. For details, see Schuler & van der Laan (2024), for example.
Theorem 4.3 (Asymptotic normality).
Suppose that Assumptions 1.1, and 4.14.2 hold. Then, the AIPW estimator converges in distribution to a normal distribution as $\sqrt{n}\Big(\widetilde{\tau}^{\mathrm{AIPW}}_{n}-\tau_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*})$, where $V^{*}$ is the efficiency bound defined as $V^{*}\coloneqq{\mathbb{E}}\left[\frac{\sigma^{2}(1,X)}{e_{0}(X)}+\frac{\sigma^{2}(0,X)}{1-e_{0}(X)}+\big(\tau_{0}(X)-\tau_{0}\big)^{2}\right]$ and $\tau_{0}(X)\coloneqq{\mathbb{E}}[Y(1)-Y(0)\mid X]$.
Here, $V^{*}$ is the efficiency bound given by the variance of the efficient influence function; that is, $V^{*}={\mathbb{E}}\left[\Psi^{*}(X,D,Y)^{2}\right]$ holds (van der Vaart, 1998). Thus, this estimator is efficient.
5 Generalization via the Bregman divergence minimization
We can further generalize our direct bias-correction term estimation from the viewpoint of Bregman divergence minimization and point out the connection to Riesz regression and covariate balancing.
5.1 Riesz regression and covariate balancing
Our resulting objective function is the same as the one used in Riesz regression (Chernozhukov et al., 2022a, 2024). Although the motivation and derivation differ, we point out that Riesz regression can be interpreted as a specific instance of DRE.
Building on this perspective, we generalize the objective function using Bregman divergence, following Sugiyama et al. (2011). Bregman divergence is a measure of discrepancy between two points, defined in terms of a strictly convex function (Bregman, 1967). In the context of DRE, Sugiyama et al. (2011) demonstrates that various existing methods can be formulated as Bregman divergence minimization problems. Inspired by this idea, we also extend the direct estimation method for the bias-correction term within the Bregman divergence minimization framework.
This generalization further allows us to derive the empirical balancing method proposed in Chan et al. (2015), which aims to achieve covariate balance and known as an instance of tailored loss function (Zhao, 2019). These results suggest that Riesz regression and covariate balancing can be unified under the DRE framework.
5.2 Bregman divergence minimization
Let $\mathcal{G}$} be a set of functions $g\colon{\mathbb{R}}\to{\mathbb{R}}$ that is differentiable and strictly convex. Given $d\in\{1,0\}$, we define the Bregman divergence between $r(d,\cdot),r(d,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ as
$\mathrm{br}^{\dagger}_{g}\big(r_{0}(d,x)\mid r(d,x)\big)\coloneqq g(r_{0}(d,x))-g(r(d,x))-\partial g(r(d,x))(r_{0}(d,x)-r(d,x)),$
where $\partial g$ denotes the derivative of $g$. Then, we define the average Bregman divergence as
$\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big].$
We estimate $r_{0}$ by $r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$. By dropping the term that is irrelevant to learning, we have
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$. For the derivation, see Appendix B.
Then, we estimate the bias-correction term $h_{0}$ (or $r_{0}$) by minimizing an empirical Bregman divergence as
$\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g}\big(r\big)+\lambda J(r),$
where $\widehat{\mathrm{BR}}_{g}(r)\coloneqq\sum_{d\in\{1,0\}}\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=d]\Big(-g(r(d,X_{i}))+\partial g(r(d,X_{i}))r(d,X_{i})\Big)-\partial g(r(d,X_{i}))\right).$
5.3 Least squares
Our least squares method for direct bias-correction term estimation can be obtained by using $g(r)=r^{2}$. Under this choice of $g$, we obtain $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\left[-2r(d,X)+\mathbbm{1}[D=d]r(d,X)^{2}\right]$. This gives an objective that is the same as the one used in Chernozhukov et al. (2022a).
5.4 Constrained maximum likelihood estimation
By using different choices of $g$, we can derive various objective functions for direct bias-correction term estimation. As a specific case of Bregman divergence minimization, we introduce constrained maximum likelihood estimation.
Consider $g^{\mathrm{L}}(r)=r\log r-r$, which is a convex function. By substituting this function into the Bregman divergence, we obtain
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big).$ (1)
where $\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big)\coloneqq{\mathbb{E}}\big[-\log(r(1,X))-\log(r(0,X))+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big]$.
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}\big(r\big)+\lambda J(r)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(-\log\big(r(1,X_{i})\big)-\log\big(r(0,X_{i})\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Sugiyama et al., 2011), which generalizes the KL importance estimation procedure (KLIEP).
Note that solving (1) is equal to
$r^{*}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\log r(d,X)\right]\quad\mathrm{s.t.}~~~{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X_{i})\big]={\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X_{i})\big]=1.$
This technique is known as Silverman’s trick (Silverman, 1982). For details, see Theorem 3.3 in Kato et al. (2023). We can replace the expected values with the sample means and define the estimation problem as $\widehat{r}_{n}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\log r(d,X_{i})\quad\mathrm{s.t.}~~~\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]r(1,X_{i})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=0]r(0,X_{i})=1$.
5.5 Empirical balancing
Next, we derive empirical balancing as a special case of Bregman divergence minimization (Chan et al., 2015). Empirical balancing is known as a specific form of covariate balancing and can be derived from a tailored loss function (Zhao, 2019).
Let us consider $g^{\mathrm{E}}(r)=(r-1)\log\left(r-1\right)-r$, which is also convex for $r\in(1,\infty)$ and its derivative is $\partial g^{\mathrm{E}}(r)=\log\left(r-1\right)$. By substituting this function, we obtain $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(\log\left(r-1\right)+r(d,X)\Big)-\log\left(r-1\right)\Big]$. Note that it holds that $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)=$
${\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(r(1,X)-1\right)-\mathbbm{1}[D=1]\log\left(r(0,X)-1\right)+\mathbbm{1}[D=1]r(1,X)+\mathbbm{1}[D=0]r(0,X)\Big].$
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}\big(r\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(r(1,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]\log\big(r(0,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. If we model $r$ as $r(1,x)=1/e(x)$ and r(0,x)=1/(1e(x))r(0,x)=1/(1-e(x)), we have r(1,x)1=1/(r(0,x)1)r(1,x)-1=1/(r(0,x)-1). Therefore, we have $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=$
$\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=1]\left(-\log\left(\frac{1}{r(1,X_{i})-1}\right)+r(1,X_{i})\right)+\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r(0,X_{i})-1}\right)+r(0,X_{i})\right)\right).$
This objective function is equivalent to the one with the tailored loss proposed in Zhao (2019). From this objective, we can also derive empirical balancing (Chan et al., 2015).
6 Discussion and related work
6.1 Related work
The estimation of the bias-correction term or the propensity score has been a core interest in causal inference. The bias-correction term can be interpreted as a gradient in a one-step estimator (van der Vaart, 2002). This idea was refined by van der Laan (2006) as targeted maximum likelihood estimation (TMLE). Chernozhukov et al. (2024) also explore a related topic from the viewpoint of Riesz representers and double machine learning (DML) (Chernozhukov et al., 2018). They propose automatic DML along with its implementation (Chernozhukov et al., 2022a).
This topic has been addressed from various perspectives. One such approach focuses on estimating the propensity score by matching the distributions of the treatment and control groups (Chan et al., 2015; Deville & Särndal, 1992; Graham et al., 2012; Bryan S. Graham & Egel, 2016; Hellerstein & Imbens, 1999). For example, Imai & Ratkovic (2013) introduces the Covariate Balancing Propensity Score (CBPS), which estimates propensity scores by explicitly balancing covariate means. In parallel with, or subsequent to, Imai & Ratkovic (2013), several methods related to covariate balancing have been proposed (Hainmueller, 2012; Zubizarreta, 2015; Athey et al., 2018).
From a methodological perspective, our study is inspired by DRE (Sugiyama et al., 2012). We review the literature of DRE in Appendix A.
Table 1: Experimental results. We report the empirical MSE and Bias of each method.
Data Dimension DM DBC (LS) DBC (UKL) Logistic CBPS RieszNet DM
Three-layer perceptron Dragonnet Linear model
IPW DR IPW DR IPW DR IPW DR IPW DM DR
Model 1 $K=3$ MSE 0.006 0.392 0.005 0.374 0.005 0.330 0.004 1.429 0.006 0.017 0.021 0.040 2.781
$K=3$ Bias -0.037 -0.299 -0.024 -0.316 -0.023 -0.257 -0.022 -0.747 -0.037 -0.027 -0.025 -0.053 -0.197
$K=3$ MSE 0.521 1.956 0.481 2.779 0.478 6.510 0.507 3.570 0.515 0.464 0.510 0.379 7.511
$K=10$ Bias 0.094 -0.930 0.086 -0.822 0.088 -0.268 0.091 -1.422 0.089 -0.093 -0.106 -0.017 0.101
Model 2 $K=3$ MSE 0.048 0.343 0.033 0.819 0.037 2.838 0.045 1.848 0.044 0.030 0.034 0.051 2.866
Bias $K=3$ -0.009 -0.275 -0.011 -0.382 -0.010 -0.403 -0.011 -0.781 -0.012 -0.022 -0.020 -0.057 -0.214
$K=3$ MSE 0.517 2.006 0.474 2.980 0.477 6.517 0.507 3.816 0.512 0.407 0.446 0.424 7.482
$K=10$ Bias 0.085 -0.944 0.082 -0.823 0.085 -0.269 0.089 -1.410 0.084 -0.087 -0.096 -0.012 0.093
6.2 Comparison with the standard DRE approaches
If we follow the standard DRE approach, we may formulate the problem as the direct estimation of $r_{0}(1,X)$. For example, when using LSIF, the risk is given by ${\mathbb{E}}\big[-2r(1,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X)^{2}\big]$, which corresponds to a part of our risk: ${\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big]$. Thus, our proposed method is closely connected to LSIF. However, the standard DRE approach does not address whether it is suitable for bias-correction term estimation. In fact, we can estimate $r_{0}$ by minimizing the LSIF risk, but our proposed method adopts a different risk: the sum of ${\mathbb{E}}\big[-2r(1,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X)^{2}\big]$ and ${\mathbb{E}}\big[-2r(0,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X)^{2}\big]$, which is directly related to the bias-correction term.
7 Simulation studies
We assess the empirical performance of our method through simulation studies, evaluating ATE estimation error across a range of scenarios. We compare our approach with CBPS (Imai & Ratkovic, 2013) and RieszNet (Chernozhukov et al., 2022a). Because our least squares is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences.
We consider two different dimensions for $X$, setting $K=3$ and $K=10$, and two different outcome models. This results in a total of four experimental settings. In all cases, the true ATE is fixed at $\tau_{0}=5.0$. To generate synthetic data, we first sample covariates $X_{i}$ from a multivariate normal distribution $\mathcal{N}(0,I_{K})$, where $I_{K}$ denotes the $K\times K$ identity matrix. The propensity score is then defined as $e_{0}(X_{i})=\frac{1}{1+\exp\bigl(-h(X_{i})\bigr)}$, where $h(X_{i})=\sum_{j=1}^{3}\alpha_{j}X_{i,j}+\sum_{j=1}^{3}\beta_{j}X_{i,j}^{2}+\gamma_{1}X_{i,1}X_{i,2}+\gamma_{2}X_{i,2}X_{i,3}+\gamma_{3}X_{i,1}X_{i,3}$. The coefficients $\alpha_{j}$, $\beta_{j}$, and $\gamma_{j}$ are independently drawn from $\mathcal{N}(0,0.5)$. Given these propensity scores, the treatment assignment $D$ is sampled accordingly. The outcome is then generated under two models, referred to as Model 1 and Model 2. In Model 1, we specify $Y_{i}=\left(X^{\top}_{i}\beta\right)^{2}+1.1+\tau_{0}D_{i}+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ and $\tau_{0}=5.0$. In Model 2, the outcome is generated as $Y_{i}=X^{\top}_{i}\beta+\left(X^{\top}_{i}\beta\right)^{2}+3\sin(X_{i,1})+1.1+\tau_{0}D_{i}+\varepsilon_{i}$.
For propensity score estimation, we employ a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer). The final output layer applies a sigmoid function to ensure that the estimated propensity scores remain in $(0,1)$. We use this model for our method, logistic regression, and CBPS. For RieszNet, we adopt the DragonNet architecture proposed in Shi et al. (2019), following the original implementation by Chernozhukov et al. (2022a). For each propensity score estimation method, including ours, we compute both the IPW and AIPW estimators using the estimated scores. Additionally, we include the direct method (DM) estimator with neural networks for comparison. In each case, the expected conditional outcomes are estimated using a three-layer neural network (100 nodes per hidden layer, with ELU activation). As a baseline, we also consider the DM estimator with linear models.
The sample size is fixed at $n=3000$. As noted earlier, we evaluate two values of $K$ ($K=3$ and $K=10$) and two outcome-model specifications (Model 1 and Model 2), resulting in four experimental configurations. Each setting is repeated 500 times. We report the MSEs and biases of the resulting ATE estimates in Table 1 for $n=3000$. Overall, the results indicate that our direct bias-correction approach achieves competitive or superior estimation accuracy compared with logistic regression and CBPS, highlighting the benefits of explicitly estimating the bias-correction term in the ATE context. RieszNet tends to outperform our method, but we consider this to be partly due to differences in the regression models. While RieszNet employs DragonNet, we use a simpler implementation. We do not employ such models, as model complexity is not our primary focus. Nevertheless, we emphasize that our method outperforms most existing approaches while exhibiting comparable performance to RieszNet.
8 Conclusion
This study proposed direct bias-correction term estimation in ATE estimation. Instead of focusing on estimating the propensity score itself, our approach directly minimizes the estimation error of the bias-correction term, leveraging empirical risk minimization techniques. We demonstrated that this direct approach enhances estimation accuracy by avoiding the intermediate step of propensity score estimation. Additionally, our method was analyzed through the lens of Bregman divergence minimization, providing a generalized framework.
Appendix A Density-Ratio Estimation (DRE)
Given two probability distributions $P$ and $Q$ over a common space $\mathcal{X}$, the density ratio function is defined as
$r_{0}(x)\coloneqq\frac{p(x)}{q(x)},$
where $p(x)$ and $q(x)$ denote the density functions of $P$ and $Q$, respectively. DRE is a fundamental problem in statistical learning, with applications in importance sampling, anomaly detection, and covariate shift adaptation.
In DRE, estimating the two densities separately can magnify estimation errors, whereas directly modeling and estimating the density ratio can lead to improved accuracy. Thus, the aim of DRE is to estimate the density ratio in an end-to-end manner by directly optimizing a single objective. Various methods for DRE have been proposed (Huang et al., 2007; Gretton et al., 2009; Qin, 1998; Cheng & Chu, 2004; Nguyen et al., 2010; Kato et al., 2019), many of which can be generalized as instances of Bregman divergence minimization (Sugiyama et al., 2011; Kato & Teshima, 2021).
Let $\mathcal{R}$ be a hypothesis class for $r_{0}$, consisting of functions $r\colon\mathcal{X}\to{\mathbb{R}}$. The goal of direct DRE is to find an optimal function $r^{*}\in\mathcal{R}$ that best approximates $r_{0}$. A natural approach is to minimize the expected squared error:
${\mathbb{E}}_{P}\left[\big(r_{0}(X)-r(X)\big)^{2}\right].$
However, since $r_{0}(x)$ is unknown, direct minimization of this objective is infeasible.
Instead, we derive an equivalent formulation that does not require knowledge of $r_{0}$. Specifically, we show that minimizing the expected squared error is equivalent to minimizing the following alternative objective:
$2{\mathbb{E}}_{Q}\left[r(X)\right]+{\mathbb{E}}_{P}\left[r(X)^{2}\right].$
This transformation enables empirical risk minimization without explicit access to the true density ratio.
Furthermore, we extend this framework by providing theoretical guarantees on the estimation error using tools from empirical process theory. From the perspective of Bregman divergence minimization, we establish a generalized methodology for DRE that accommodates various estimation strategies.
Finally, we present numerical experiments that demonstrate the effectiveness of our approach in practical scenarios, including importance weighting and outlier detection.
Appendix B Bregman divergence
Here, we show the equivalence between
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big),$
where $\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$, and
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$.
This can be shown as follows:
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(r(d,x))r_{0}(d,X)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(r(d,x))\Big].$
Here, we used
${\mathbb{E}}[\mathbbm{1}[D=1]r_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)r_{0}(1,X)\mid X]=1.$
We can show the same result for the case with $D=0$.
Appendix C Estimation of the average treatment effect for the treated (ATT)
Our method can also be applied to other estimands, such as the ATT, which is defined as
$\alpha_{0}\coloneqq{\mathbb{E}}\big[Y(1)-Y(0)\mid D=1\big].$
The IPW and AIPW estimators designed for the ATT are given by
IPW estimator.
$\widetilde{\alpha}^{\mathrm{IPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]Y_{i}}{\pi_{0}}-\frac{e_{0}(X_{i})\mathbbm{1}[D_{i}=0]Y_{i}}{\pi_{0}(1-e_{0}(X_{i}))}\right)=\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))}\right)Y_{i}$.
AIPW estimator.
$\widetilde{\alpha}^{\mathrm{AIPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))}\right)\left(Y_{i}-\mu_{0}(0,X_{i})\right)$,
where $\pi_{0}={\mathbb{E}}[\mathbbm{1}[D=1]]$.
Thus, the bias-correction term for ATT estimation is given as
$\widetilde{h}_{0}(D,X,e_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
where $\pi_{0}={\mathbb{E}}[\mathbbm{1}[D=1]]$.
Let $w_{0}(x)\coloneqq\frac{e_{0}(X)\mathbbm{1}[D=0]}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
$\widetilde{h}_{0}(D,X,w_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
Let $\mathcal{W}$} be a set of functions $w\colon{\mathcal{X}}\to{\mathbb{R}}_{+}$. Then, we define the following least squares:
$w^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(\widetilde{h}(D,X;r_{0},\pi_{0})-\widetilde{h}(D,X;r,\pi_{0})\big)^{2}\right].$
Note that we use $\pi_{0}$ itself. We can show that this least squares is equivalent to
$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}_{1}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
where ${\mathbb{E}}_{1}$ is expectation over the treated group ($p(x\mid d=1)$). The empirical version of this risk is given as
$\widehat{w}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2\frac{1}{\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]w(X_{i})+\frac{1}{n}\sum^{n}_{i=1}w(X_{i})^{2}\right\},$
We can demonstrate the equivalence between the two least-squares formulations as follows:
$w^{*}$ $=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(\widetilde{h}(D,X;r_{0},\pi_{0})-\widetilde{h}(D,X;r,\pi_{0})\big)^{2}\right]$
$=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(w_{0}(X)\mathbbm{1}[D=0]-w(X)\mathbbm{1}[D=0]\big)^{2}\right]$
$=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[-2w_{0}(X)w(X)\mathbbm{1}[D=0]+w(X)^{2}\mathbbm{1}[D=0]\right].$
To see this equivalence, consider
${\mathbb{E}}\left[w_{0}(X)w(X)\mathbbm{1}[D=0]\right]$
$={\mathbb{E}}\left[{\mathbb{E}}\left[w_{0}(X)w(X)(1-e_{0}(X))\right]\right]$
$={\mathbb{E}}\left[e_{0}(X)w(X)/\pi_{0}\right]$
$=\int\frac{1}{\pi_{0}}e_{0}(x)w(x)\pi_{0}(x){\mathrm{d}}x$
$=\int\frac{1}{\pi_{0}}\frac{\pi_{0}p_{0}(x\mid d=1)}{p_{0}(x)}w(x)p_{0}(x){\mathrm{d}}x$
$=\int p_{0}(x\mid d=1)w(x){\mathrm{d}}x.$
This confirms the equivalence between the two least-squares objectives.
Appendix D Preliminary
This section introduces notions that are useful for the theoretical analysis.
D.1 Rademacher complexity
Let $\sigma_{1},\dots,\sigma_{n}$ be $n$ independent Rademacher random variables; that is, independent random variables for which $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$. Let us define
$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(X_{i}).$
Additionally, given a class $\mathcal{F}$}, we define
$\mathfrak{R}_{n}{\mathcal{F}}\coloneqq\sup_{f\in{\mathcal{F}}}\mathfrak{R}_{n}f.$
Then, we define the Rademacher average as ${\mathbb{E}}[\mathfrak{R}_{n}{\mathcal{F}}]$ and the empirical Rademacher average as ${\mathbb{E}}_{\sigma}[\mathfrak{R}_{n}{\mathcal{F}}\mid X_{1},\dots,X_{n}]$.
D.2 Local Rademacher complexity bound
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. For $f\in{\mathcal{F}}$, let us define
$Pf\coloneqq{\mathbb{E}}[f(X)],$
$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}).$
We introduce the following result about the Rademacher complexity.
Proposition D.1 (From Theorem 2.1 in Bartlett et al. (2005)).
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(X))\leq r$. Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
$\sup_{f\in{\mathcal{F}}}\Big(Pf-P_{n}f\Big)\leq\inf_{\alpha>0}\left\{2(1+\alpha){\mathbb{E}}[\mathfrak{R}_{n}f]+\sqrt{\frac{2rx}{n}}+(b-a)\left(\frac{1}{3}+\frac{1}{\alpha}\right)\frac{z}{n}\right\}.$
D.3 Bracketing entropy
We define the bracketing entropy. For a more detailed definition, see Definition 2.2 in van de Geer (2000).
Definition D.2.
Bracketing entropy. Given a class of functions $\mathcal{F}$}, the logarithm of the smallest number of balls in a norm $\|\cdot\|_{2,P}$ of radius $\delta>0$ needed to cover $\mathcal{F}$} is called the $\delta$-entropy with bracketing of $\mathcal{F}$} under the $L_{2}(P)$ metric, denoted by $H_{B}(\delta,{\mathcal{F}},P)$.
D.4 Talagrand’s concentration inequality
We introduce Talagrand’s lemma.
Proposition D.3 (Talagrand’s Lemma).
Let $\phi\colon{\mathbb{R}}\to{\mathbb{R}}$ be a Lipschitz continuous function with a Lipschitz constant $L>0$. Then, it holds that
$\mathfrak{R}_{n}(\phi\circ{\mathcal{F}})\leq L\mathfrak{R}_{n}({\mathcal{F}}).$
Appendix E Basic inequalities
Recall that we have defined an estimator $\widehat{r}$ as follows:
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where $J(r)$ is some regularization term.
Throughout the proof, we use the following basic inequalities that hold for $\widehat{r}$.
Proposition E.1.
The estimator $\widehat{r}$ satisfies the following inequality:
[…]
$\leq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r_{0}(1,X_{i})-2r_{0}(0,X_{i})+\mathbbm{1}[D_{i}=1]r_{0}(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r_{0}(0,X_{i})^{2}\Big)+\lambda J(r_{0}).$
For a function $f\colon{\mathcal{X}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean and expectation as
$\int f{\mathrm{d}}P_{n}$ $\coloneqq\int f(\cdot){\mathrm{d}}P_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}),$
$\int f{\mathrm{d}}P$ $\coloneqq\int f(\cdot){\mathrm{d}}P\coloneqq{\mathbb{E}}[f(X)].$
Following this notation, we also express this inequality as
$\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int\mathbbm{1}[D_{i}=1]\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int\mathbbm{1}[D_{i}=0]\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,X)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\lambda J(\widehat{r})-\lambda J(r_{0})\leq 0.$
Proposition E.2.
The estimator $\widehat{r}$ satisfies the following inequality:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid X)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
[…]
Proof of Proposition E.1 is trivial. We prove Proposition E.2 below.
Proof.
Let $e_{0}(1\mid x)=e_{0}(x)$ and $e_{0}(0\mid x)=1-e_{0}(x)$. Then, the following holds for the $L_{2}$ estimation error:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$=\left\|\mathbbm{1}[D=1]\big(\widehat{r}(1,X)-r_{0}(1,X)\big)-\mathbbm{1}[D=0]\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\left\|e_{0}(1\mid\cdot)\big(\widehat{r}(1,X)-r_{0}(1,X)\big)\right\|^{2}_{P_{0}}+\left\|e_{0}(1\mid 0)\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}.$
From Proposition E.1, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0})$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\int-2\big(r_{0}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(r_{0}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int e_{0}(0,\cdot)\Big(r_{0}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0}).$
Here, we used
$\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$ $=2\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$
$=2\int r_{0}(1,\cdot){\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}.$
Appendix F Proof of Theorem 3.5
We show Theorem 3.5 by bounding
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$ (2)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$ (3)
in Proposition E.2.
Here, since $r_{0}\in{\mathcal{H}}$, it holds that $r^{*}=r_{0}$, which implies that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)=0.$
Therefore, we consider bounding the first sum
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg).$
We can bound this term by using the empirical-process arguments.
F.1 Preliminary
The following proposition is originally presented in van de Geer (2000) and was rephrased by Kanamori et al. (2012) in a form that is convenient for the theoretical analysis in DRE.
Lemma F.1 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
Let $\mathcal{F}\subset L^{2}(P)$ be a function class and the map $I(f)$ be a complexity measure of $f\in\mathcal{F}$, where $I$ is a non-negative function on $\mathcal{F}$ and $I(f_{0})<\infty$ for a fixed $f_{0}\in\mathcal{F}$. We now define $\mathcal{F}_{M}=\{f\in\mathcal{F}:I(f)\leq M\}$ satisfying $\mathcal{F}=\bigcup_{M\geq 1}\mathcal{F}_{M}$. Suppose that there exist $c_{0}>0$ and $0<\gamma<2$ such that
$\sup_{f\in\mathcal{F}_{M}}\|f-f_{0}\|\leq c_{0}M,\ \sup_{\stackrel{{\scriptstyle f\in\mathcal{F}_{M}}}{{\|f-f_{0}\|_{L^{2}(P)}\leq\delta}}}\|f-f_{0}\|_{\infty}\leq c_{0}M,\quad\text{for all }\delta>0,$
and that $H_{B}(\delta,\mathcal{F}_{M},P)=O{M/\delta}^{\gamma}$. Then, we have
$\sup_{f\in\mathcal{F}}\frac{\left|\int(f-f_{0})d(P-P_{n})\right|}{D(f)}=O_{p}(1),\ (n\to\infty),$
where $D(f)$ is defined by
$D(f)=\max{\frac{\|f-f_{0}\|_{L^{2}(P)}^{1-\gamma/2}I(f)^{\gamma/2}}{\sqrt{n}}}{\frac{I(f)}{n^{2/(2+\gamma)}}}.$
F.2 Upper bound using the empirical-process arguments
Proposition F.2.
[From Lemma 2 in Kanamori et al. (2012)] Under the conditions of Theorem 3.5, for any $0<\gamma<2$, we have
$\left|\int\left(\widehat{r}(d,\cdot)-r^{*}(d,\cdot))({\mathrm{d}}(P_{0}-P_{n})\right)\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left\|\widehat{r}(d,\cdot)\right\|^{\gamma/2}_{{\mathcal{H}}}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
$\left|\int\left(\widehat{r}^{2}(d,\cdot)-r^{*}(d,\cdot)^{2}\right)({\mathrm{d}}(P_{0}-P_{n}))\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|^{2}_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
as $n\to\infty$.
F.3 Proof of Theorem 3.5
We prove Theorem 3.5 following the arguments in Kanamori et al. (2012).
Proof.
From Proposition E.2 and $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$\leq\sum_{d\in\{1,0\}}\left|\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})\right|+\left|\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\right|+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
From Proposition F.2, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$=\sum_{d\in\{1,0\}}O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
We consider the following three possibilities:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (4)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (5)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (6)
The above inequalities are analyzed as follows:
Case (4).
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=O_{p}(\lambda),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
Therefore, we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
Case (5).
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
From the first inequality, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
By using this result, from the second inequality, we have
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\left(\frac{1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
This implies that
$\|\widehat{r}\|_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
Therefore, the following inequity is obtained.
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Case 6.
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
As well as the argument in (5), we have $\|\widehat{r}\|_{{\mathcal{H}}}=o_{p}(1)$. Therefore, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Appendix G Proof of Theorem 3.8
Our proof procedure mainly follows those in Kato & Teshima (2021) and Zheng et al. (2022). In particular, we are inspired by the proof in Zheng et al. (2022).
We prove Theorem 3.8 by proving the following lemma:
Lemma G.1.
Suppose that Assumption 3.7 holds. For any $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, there exists a constant $C>0$ depending on $(\mu,\sigma,M)$ such that for any $\gamma>0$, with probability at least $1-\exp(-\gamma)$, it holds that
$\Big\|\widehat{f}_{n}-f_{0}\Big\|_{2}\leq C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\big\|f^{*}-f_{0}\big\|_{2}+\sqrt{\frac{\gamma}{n}}\right).$
As shown in Zheng et al. (2022), we can bound $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)$ by specifying neural networks and obtain Theorem 3.8.
G.1 Proof of Lemma G.1
We prove Lemma G.1 by bounding (2) in Proposition E.2.
To bound (2), we show several auxiliary results. Define
$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(X_{i})-f^{*}(X_{i}))^{2}\leq u\},$
$\overline{{\mathcal{G}}}^{f^{*},u}$ $\coloneqq\Big\{(f-f^{*})\colon f\in\widehat{{\mathcal{F}}}^{f^{*},u}\Big\},$
$\kappa^{u}_{n}(u)$ $\coloneqq{\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u}\right],$
$u^{\dagger}$ $\coloneqq\inf\big\{u\geq 0\colon\kappa^{u}_{n}(s)\leq s^{2}\ \ \ \forall s\geq u\big\}.$
Here, we show the following two lemmas:
Lemma G.2 (Corresponding to (26) in Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,x)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq C\left(\|f^{*}(X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(X)-f_{0}(X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
Lemma G.3 (Corresponding to (29) in Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
Additionally, we use the following three propositions directly from Zheng et al. (2022).
Proposition G.4 (From (32) in Zheng et al. (2022)).
Let $u>0$ be a positive value such that
$\|f-f_{0}\|_{2}\leq u$
for all $f\in{\mathcal{F}}$. Then, for every $z>0$, with probability at least $1-2\exp(-z)$, it holds that
$\sqrt{\frac{1}{n}\sum^{n}_{i=1}\big(f(X_{i})-f_{0}(X_{i})\big)^{2}}\leq 2u.$
Proposition G.5 (Corresponding to (36) in Step 3 of Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. Then, there exists a universal constant $C>0$ such that
$u^{\dagger}\leq CM\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}.$
Proposition G.6 (Upper bound of the Rademacher complexity).
Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $r_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
Then, we prove Lemma G.1 as follows:
Proof of Lemma G.1.
If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then from (2) and Lemmas G.2 and G.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq C\left(\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}+u_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (7)
This result implies that if $\sqrt{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})}$, then there exists $n_{0}$ such that for all $n>n_{0}$, there exists $u_{1}<u_{0}$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}\leq u_{1}.$
For any $z>0$, define $\overline{u}$ as
$\overline{u}_{z}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}.$
Define a subspace of ${\mathcal{F}}^{\mathrm{FNN}}$ as
${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}_{z}\coloneqq\big\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\|f-f_{0}\|\leq\overline{u}_{z}\big\}.$
Define
$\ell\coloneqq\lfloor\log_{2}(2M/\sqrt{\log(n)/n})\rfloor.$
Using the definition of subspaces, we divide ${\mathcal{F}}^{\mathrm{FNN}}$ into the following $\ell+1$ subspaces:
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{0}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}),$
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{1}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u})\backslash{\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}),$
$\quad\vdots$
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{\ell}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},2^{\ell}\overline{u})\backslash{\mathcal{S}}^{\mathrm{FNN}}(f_{0},2^{\ell-1}\overline{u}).$
Since $\overline{u}_{z}>u^{\dagger}$, from the definition of $u^{\dagger}$, we have
$\overline{u}^{2}_{z}\leq\kappa^{u}_{n}(\overline{u}).$
If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{j}$, then from (G.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{2}$
$\leq C\left(2^{\ell-1}\overline{u}\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)+\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (8)
Additionally, if
$C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)\leq\frac{1}{8}2^{j}\overline{u},$ (9)
$C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right)\leq\frac{1}{8}2^{2j}\overline{u}^{2}$ (10)
hold, then
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{2}\leq 2^{j-1}\overline{u}.$ (11)
Here, to obtain (11), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (G.1), (9), and (10).
From Proposition G.5, it holds that
$u^{\dagger}\leq CM\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}.$
Therefore, we can choose $\overline{u}$ as
$\overline{u}\coloneqq C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\log(n)/n}+4\sqrt{3}M\sqrt{z/n}\right),$
where $C>0$ is a constant independent of $n$. ∎
G.2 Proof of Lemma G.2
From Proposition D.1, we have
$\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{1}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n},$
$\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{2}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}.$
This is a direct consequence of Proposition D.1. Note that $r^{*}$ and $r_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition D.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part. Here, we also used
$\Big|r^{*}(d,x)-r_{0}(d,x)\Big|\leq\frac{1}{4}\Big|f^{*}(x)-f_{0}(x)\Big|,$
$\Big|r^{*}(d,x)^{2}-r_{0}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f^{*}(x)-f_{0}(x)\Big|.$
We have
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{0}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}\Bigg)$
$\leq C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16CMz}{3n}\right).$
G.3 Proof of Lemma G.3
Let $g\coloneqq(f-f^{*})^{2}$. From the definition of FNNs, we have
$g\leq 4M^{2}$
Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq r_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}r^{2}$.
Here, we note that the followings hold for all $f$ ($r$):
$\Big|r(d,x)-r^{*}(d,x)\Big|\leq\frac{1}{4}\Big|f(x)-f^{*}(x)\Big|,$
$\Big|r(d,x)^{2}-r^{*}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f(x)-f^{*}(x)\Big|.$
Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that, for each $d\in\{1,0\}$ it holds that Then, it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+r_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$